ashten0219:

The tickets in a box are numbered from 1 to 15 inclusive. If a ticket is drawn at random and replaced and then a second is drawn at random, what is the probability that the sum of their numbers is even? A: 49/225 B: 1/2 C: 113/225

ashten0219:

I have no idea how or what to do to solve this.

Shadow:

Changed the subject to mathematics. Be sure you're posting these in the right section.

ashten0219:

How do I ensure that?

Shadow:

For example, if someone is tabbed into the Mathematics section, they wouldn't see this on the question list and be able to help you, as you're putting it in Misc.

Shadow:

When you begin typing into the 'Ask a question' box in the top left, there is a white box under the bottom left corner which allows you to choose a subject. Select Mathematics/the appropriate subject.

ashten0219:

Did it change now?

Shadow:

This question is a bit more abstract, I might have to actually turn on my brain.

Shadow:

And I already put it in Mathematics.

ashten0219:

Oh, thank you. I'll remember that in the future.

Shadow:

So we have: 1, 2, 3, 4, 5, 6,7, 8, 9, 10, 11, 12, 13, 14, 15 odds: 1, 3, 5, 7, 9, 11, 13, 15 evens: 2, 4, 6, 8, 10, 12, 14 We have 8 odds and 7 evens.

Shadow:

So when you get: odd + odd, the sum is an even number odd + even, the sum is an odd number even + even, the sum is an even number

ashten0219:

How are there 225 possibilities?

Shadow:

Since there's two choices and the first choice ticket gets 'replaced' it is 15 outcomes times 15 outcomes, 15^2, which is 225.

Shadow:

If the first ticket didn't get replaced, it'd be 15 times 14.

ashten0219:

so since there are more odd numbers, would that make it more likely that the sum will be odd?

ashten0219:

since the odd and even make an odd

ashten0219:

or would it be more likely that it'll be even because an odd and an odd make an even

Shadow:

Well when you draw a ticket there's a 8/15 chance it's odd, and a 7/15 chance it's even. So slightly more of the time it will be odd, and it maintains that slight advantage since I assume it get's replaced with the same number it was. And as I posted, odd + odd, the sum is an even number odd + even, the sum is an odd number even + even, the sum is an even number

ashten0219:

so would I multiply 8/15 and 7/15?

Shadow:

So 7 times 7 presents the 49 outcomes where two even #s were chosen

Shadow:

And 8 times 8 presents the outcomes where it's odd + odd

ashten0219:

Oh so we only multiply the evens?

Shadow:

We multiply the outcomes where the sum is even

Shadow:

Remember odd + odd is even

Shadow:

1 + 3 = 4 5 + 7 = 12

Shadow:

So all outcomes for odd + odd which is 8 times 8, and all outcomes of even + even, which is 7 times 7

Shadow:

Do the math, add the sum of both outcome sets, and that should be your # of outcomes where the sum of the two tickets chosen is even.

ashten0219:

so 7 times 7 = 49 15 times 15 = 225 so the answer should be 49/225?

Shadow:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shadow So all outcomes for odd + odd which is 8 times 8, and all outcomes of even + even, which is 7 times 7 \(\color{#0cbb34}{\text{End of Quote}}\)

ashten0219:

I'm confused lol

Shadow:

So you want the probability that the sum of the two tickets chosen is even

Shadow:

So you determine which outcomes make the sum even

Shadow:

Which is odd + odd and even + even

Shadow:

Since odd + even makes an odd #, for example: 1 + 4 = 5, 5 + 4 = 9

ashten0219:

so multiply 8 times 8 and 7 times 7 and add them together?

Shadow:

Then you take all the outcomes where the tickets are odd + odd, and even + even. Two odds is 8 times 8 Two evens is 7 times 7

Shadow:

Since there are 8 odds and 7 evens

ashten0219:

I'm really sorry math is my absolute worst subject and I'm slow as it is. so multiply 8 and 8 and 7 and 7 add them together and then 15 times 15

Shadow:

Yes with the first part being the numerator (outcomes where the sum of the ticket numbers is even) and the 15 times 15 being the denominator, or total outcomes/sample size.

Shadow:

And you're fine man

ashten0219:

8*8= 64 7*7 = 49 64+49 = 113 and 15*15= 225 so the answer is 113/225?

Shadow:

Yes

ashten0219:

omgosh thank you!

Shadow:

yw